# Magnetic Dipole Portal Vector Dark Matter at Fixed-Targets

### v1

@@ 4.2 Direct Detection
Here, we adopt the notation of Ref. Catena:2019gfa where $m_{e}$ is the electron mass, $\boldsymbol{q}$ is the momentum transfer, $\boldsymbol{v}^{\perp}$ is the transverse DM-electron relative velocity while $\boldsymbol{S}_{e}^{s^{\prime}s}=\xi^{\dagger s^{\prime}}\boldsymbol{S}_{e}\xi^{s}$, $\boldsymbol{S}_{e}=\boldsymbol{\sigma}/2$ is the electron spin operator, $\boldsymbol{\sigma}=(\sigma_{1},\sigma_{2},\sigma_{3})$ is a three-dimensional vector whose components are the three Pauli matrices, $\xi^{s}$, $s=1,2$, is a two-component spinor, $\boldsymbol{S}_{DM}^{\lambda^{\prime}\lambda}=-i\boldsymbol{e}^{\prime}_{\lambda^{\prime}}\times\boldsymbol{e}_{\lambda}$, $\mathcal{S}^{\lambda^{\prime}\lambda}_{{\rm DM}\,ij}=\frac{1}{2}\left(e_{\lambda i}\,e^{\prime}_{\lambda^{\prime}j}+e_{\lambda j}\,e^{\prime}_{\lambda^{\prime}i}\right)$ and $e_{\lambda i}$ is the $i$-th component of the $\lambda$-th DM polarization vector. From Eq. LABEL:eq:MNR and the null results reported by DAMIC-M DAMIC-M:2025luv and PANDAX-4T Zhang:2025ajc , we derive 90% C.L. exclusion limits in the $(m_{\rm DM},\sin\zeta)$ plane. For DAMIC-M, we perform this calculation by requiring that the probability of observing a total number of signal plus background events smaller than or equal to the actually observed one in the one, two and three electron channels is at least 0.1. In doing so, we use the expected background reported in Tab. 1 of DAMIC-M:2025luv , and compute the rate of DM-induced electronic excitations in a silicon detector using the crystal response functions obtained in Catena:2021qsr . We assume the total detector efficiency given in DAMIC-M:2025luv and an exposure of 1.257 kg-year. For PANDAX-4T, we impose that, in each of the eight ionization-charge bins considered in Ref. Zhang:2025ajc , the probability of observing a number of signal plus background events smaller than or equal to the actually observed one is at least 0.1. We extract the PANDAX-4T background model from Fig. 2 of Ref. Zhang:2025ajc and compute the differential electron recoil energy rate using the xenon response functions derived in Ref. Catena:2019gfa . We then convert this into an ionization rate by applying the detector response model introduced in Ref. Essig:2012yx , the total detector efficiency reported in Fig. 1 of Ref. Zhang:2025ajc and employing an exposure of 1.04 ton-year. For both experiments, we assume Poisson statistics for the observed number of counts.
4.3 Electroweak Precision Measurements
New physics which affect SM electroweak precision observables are constrained. Following Burgess:1993vc and Alonso-Gonzalez:2025xqg , we check that precision measurements of the mass of the Z-boson and the coupling between the Z-boson and SM fermions do not exceed current values ParticleDataGroup:2024cfk . We find that these limits are not competitive with other limits considered in this study, but we leave a more comprehensive analysis of contributions from the extended scalar sector to future work. Furthermore, the invisible branching ratio of Higgs decays are excluded above 0.145 at 95% confidence level ATLAS:2022yvh . In this scenario, the process $H\to Z^{\prime}Z^{\prime}$ is found to be small, with a value of $\text{BR}_{H\to Z^{\prime}Z^{\prime}}\approx 4.6\times 10^{-13}$ for $\sin\zeta=0.1$ and $m_{\text{DM}}=1$ GeV.
4.4 Astrophysical and Cosmological Constraints
Energy Injection into the CMB
The annihilation of DM particles in the early universe injects energy into the CMB, which is constrained by current observations since it can change the recombination history, thus modifying the measured temperature and polarization power spectra Planck:2015fie . For constraints from the CMB to be relevant, DM must annihilate with a cross section that is s-wave, such that it is not velocity suppressed [^3] There does indeed exist bounds on p-wave annihilating DM Liu:2016cnk , however they impose constraints that are not competitive with the other constraints we already consider in this work.. In this model, DM annihilation into photons through the diagrams displayed in Fig. 2 are s-wave dominant, therefore the cross section is constant at low temperatures. We thus apply the 95% confidence level bound derived in Slatyer:2015jla on the channel $\text{DM DM}\to\gamma\gamma$.
Big Bang Nucleosynthesis
Dark sector particles, especially of sub-GeV mass, thermally coupled to the SM and present during Big Bang Nucleosynthesis (BBN) can lead to alterations in the number of relativistic degrees of freedom, $N_{eff}$, affecting primordial abundances of light elements. The abundances of light elements are measured to a given precision, and any expected deviations from these observations lead to constraints on dark sector models Giovanetti:2021izc ; Sabti:2019mhn ; Depta:2019lbe . A detailed consideration of constraints from BBN is beyond the scope of this work, however we include a conservative lower bound on the mass of the lightest dark species, $Z^{\prime}$, to be >10>10 MeV. Thanks to the large dark coupling values, $g_{D}=10^{-4},10^{-3},10^{-2}$, the $Z^{\prime}$ remains in thermal equilibrium with the SM thermal bath for all values of $\sin\zeta$ (down to $\sin\zeta=10^{-10}$), therefore $m_{Z^{\prime}}>10$ MeV applies across the entire parameter space considered here.
Indirect Detection
Searches of DM annihilation products in astrophysical environments constrain the annihilation cross section of DM Cirelli:2025rky . DM annihilation into SM fermion final states such as $e^{+}e^{-}$, $q\bar{q}$, proceed via p-wave (velocity) suppressed cross sections, leaving constraints through these channels weak. We thus only consider the velocity independent annihilation into $\gamma\gamma$ channel, drawn in Fig. 2 (right), where the dominant constraints come from CMB as discussed above.
Bullet Cluster
DM self-interactions are constrained by measurements of the colliding galaxy cluster, the Bullet Cluster Markevitch:2003at ; Wittman:2017gxn . The cross section of DM self-interactions at non-relativistic energies, for all parameter values considered in this study, fall safely below the upper limit of $\sigma/m_{\text{DM}}<1\>\text{cm}/\text{g}$ reported in Markevitch:2003at .
@@ 5 Sensitivity at Fixed-Target Experiments
In this section, we consider the most relevant experimental constraints and projections, including constraints from fixed target experiments, colliders, direct detection, electroweak precision measurements, BBN, and from energy injection into the CMB.
@@
With the ongoing development of future fixed target experiments, such as LDMX LDMX:2025pkp ; LDMX:2018cma , and the current competitive limits coming from NA64 NA64:2025ddk and electron beam dump experiments Andreas:2012mt , fixed target experiments are capable of extensively probing sub-GeV DM for numerous theoretical scenarios Catena:2023use ; Catena:2025fsl ; Berlin:2020uwy ; Krnjaic:2022ozp ; Gori:2022vri ; Cosme:2021baj . Having a high energy beam of electrons (or protons in the case of SHiP SHiP:2020noy and other proton beam dumps) incident on a stationary target dense with nuclei, allows for a large number of potential DM events that can be detected through their missing energy or momentum (or through the direct scattering with a downstream detector in the case of proton beam dumps). However, for our scenario since $m_{Z^{\prime}}<m_{\text{DM}}$, production of DM proceeds through off-shell processes and is thus suppressed relative to its on-shell counterpart. The on-shell production of visible species, in particular $e^{+}e^{-}$ also occurs, acting as a complementary probe of this model. We consider the missing momentum search constraints from current NA64 null results and the sensitivity projections of the future experiment LDMX, in addition to the visible searches from past beam dump experiments E774 Bross:1989mp , E141 Riordan:1987aw , Orsay Davier:1989wz , E137 Bjorken:1988as , CHARM Gninenko:2012eq , KEK Konaka:1986cb , and NuCal Tsai:2019buq .
@@
NA64’s 100 GeV electron beam incident on an active target composed of lead absorber layers, with $4.4\times 10^{11}$ electrons on target (EOT) so far, places competitive constraints on DM and dark sector particles that can be produced through these beam target interactions NA64:2023wbi ; NA64:2025ddk . We consider both visible and invisible production, where for the latter the production of DM through dark bremsstrahlung and through the decay of vector mesons are each capable of delivering missing energy signatures.
@@
LDMX is a future missing-momentum experiment that will operate in two phases, where the second phase plans to have $10^{16}$ EOT with a beam energy of 8 GeV LDMX:2023zbn ; LDMX:2018cma . With the ability to measure the transverse momentum, in addition to the energy, of the recoil electron after the interaction with the tungsten target, LDMX is subject to less background, projecting sensitivities reaching further than previous experiments in the on-shell case. In this study we explore the less studied and less sensitive off-shell / visible decay regime Berlin:2018bsc . Similarly to NA64, LDMX will search for signatures from visible decays, dark bremsstrahlung, and invisible vector meson decay Schuster:2021mlr .
@@ 5.1 Dark Bremsstrahlung
Beam electrons incident on a target could produce dark sector particles in a process similar to ordinary bremsstrahlung. In the usual dark photon kinetic mixing scenario where $m_{A^{\prime}}>2m_{\text{DM}}$ Berlin:2020uwy , the dark photon is produced on-shell and later decays to DM, therefore the event rate is independent of the DM properties and is only dependent on the dark photon. In our scenario however, there is the off-shell production of DM through a virtual $Z^{\prime}$ or ordinary photon, since the mass hierarchy is reversed such that $m_{\text{DM}}>m_{Z^{\prime}}$. Off-shell production of DM is suppressed compared to on-shell production Berlin:2018bsc , leading to less expected events. Dark bremsstrahlung processes that occur dominantly for our choice of model parameters, namely the emission of DM off the initial or final state electron, and dark Bethe-Heitler trident processes, are drawn in Fig. 4.
Figure 4: The dominant off-shell dark bremsstrahlung processes generating DM that occur at fixed target experiments, where NN represents the target nucleus. Left: DM production diagrams off the initial or final state beam electron. Right: Bethe-Heitler trident diagrams for DM production.
Dark bremsstrahlung diagrams involving the new scalars, $\Phi_{D}^{0}$, $\Sigma_{D}^{0}$, and $\Sigma_{D}^{\pm}$, as intermediate states, become relevant as $\sin\beta$ approaches 0, however for the choice of parameters where $\cos\beta=0.7$, these diagrams contribute negligibly to the production of dark states. Simulations of dark bremsstrahlung events are performed in MadGraph Alwall:2014hca , with UFO files Degrande:2011ua generated by our FeynRules model file Alloul:2013bka . The DM and dark scalar vertices, in addition to the nucleus-photon vertex with form factor from Bjorken:2009mm  [^4] The form factor from A18 and A19 of Bjorken:2009mm , which is used as an approximate representation of the effects from the atomic/nuclear structure, has a typo in Eq. A19: the second term should not be squared., are included in our UFO files.
@@ 5.2 Dark Higgs-strahlung
Not only can the production of DM itself produce a missing momentum signature, but also the production of new scalars Acanfora:2025ikj ; Duerr:2017uap , since similarly to DM, dark scalars would not interact in the detector. In our set up, the production of scalars $\Sigma_{D}^{\pm/0},\Phi_{D}^{0}$ through the process $e^{-}N\to e^{-}NS$, where $S$ is final states such as $\Sigma_{D}^{+(-)}X^{-(+)}$, $\Sigma_{D}^{+}\Sigma_{D}^{-}$, and $\Sigma_{D}^{0}/\Phi_{D}^{0}$, would look like a missing energy/momentum signature at an experiment like LDMX, similarly to a signature from DM production. The final states $\Sigma_{D}^{+(-)}X^{-(+)}$ make up the largest contribution of events. The processes with final states $\Phi_{D}^{0}Z^{\prime}$ and $\Sigma_{D}^{0}Z^{\prime}$ also occur, however there is only missing energy/momentum if the $Z^{\prime}$ decays visibly outside of the detector.
@@ 5.3 Invisible Vector Meson Decay
In addition to the dark bremsstrahlung process, DM can also be produced invisibly through vector meson decays generated in beam-target collisions – the so called invisible vector meson decay Schuster:2021mlr . Bremsstrahlung photons are converted to vector mesons through exclusive photoproduction processes, which then decay to DM through a mediator that mixes with vector mesons. The three possible vector mediators in this model are the photon, $Z$ boson, and $Z^{\prime}$. The $Z$ boson is heavy compared to the vector mesons, resulting in a much smaller cross section than those of the other two mediators. There is a larger cross section through the ordinary photon, however not as high as through the $Z^{\prime}$ – which has a mass less than the DM but similar order of magnitude. Therefore, we consider the process $V\to Z^{\prime}\to X^{+}X^{-}$, where $V$ is a vector meson such as the $\rho$, $\omega$, $\phi$, or $J/\psi$, and the $Z^{\prime}$ is off-shell. Since the $Z^{\prime}$ couples to SM quarks through vector currents, the $Z^{\prime}$ mixes with the vector mesons with the form factors given in the fourth line of Table III of Schuster:2021mlr . Following the methodology of Schuster:2021mlr , and adapting to this model, we compute the predicted number of DM events from invisible vector meson decays for the experiments NA64 NA64:2025ddk and LDMX LDMX:2025pkp ; LDMX:2018cma . The number of events from the decay of a vector meson $V$ is given by,
@@
where $N_{V}$ is the predicted number of vector mesons produced from beam target collisions and $Br(V\to X^{+}X^{-})\equiv\frac{\Gamma_{V\to X^{+}X^{-}}}{\Gamma_{V}}$ where the decay rate to DM is written in Appendix A and $\Gamma_{V}$ is measured experimentally and reported in PDG ParticleDataGroup:2018ovx . The sum runs over $V=\rho,\omega,\phi$ and in the case of NA64, additionally $J/\psi$ since the beam energy of 100 GeV is sufficiently high. We use the corresponding values found in Table II of Schuster:2021mlr for $N_{V}$, calculated using a combination of experimental measurements and theory. We take the currently reported EOT at NA64 to be $4.4\times 10^{11}$, and thus scale-up the corresponding values in Table II of Schuster:2021mlr to be $N_{\rho}=1.1\times 10^{7}$, $N_{\omega}=7.9\times 10^{5}$, $N_{\phi}=7.0\times 10^{5}$, and $N_{J/\psi}=9.7\times 10^{3}$. For LDMX, we take the corresponding values for Phase II reported in Table II of Schuster:2021mlr .
Figure 6: 90% confidence level constraints from current NA64 null results NA64:2025ddk (in solid) and projections for LDMX Phase II LDMX:2025pkp (in dashed), from invisible vector meson decay. The contribution from each vector meson is plotted in colour, and the total in black. $g_{D}$ is set to $10^{-2}$.
90% confidence level constraints and projections for invisible vector meson decays at NA64 and LDMX, respectively, for each specie of meson are plotted in Fig. 6 as a function of the model parameters. The reach of LDMX surpasses that of NA64 due to the increased statistics, providing a significant improvement to the sensitivity landscape from the data that will come from LDMX. Interestingly, since the $V\to X^{+}X^{-}$ decay rate is proportional to $1/m_{\text{DM}}^{4}$, the sensitivity increases as $m_{\text{DM}}$ decreases – resulting in much more sensitivity for lower masses compared to the on-shell dark photon studied in Schuster:2021mlr , which features a sensitivity that is independent of $m_{\text{DM}}$. Due to the inverse mass hierarchy, where $m_{\text{DM}}>m_{Z^{\prime}}$, the sensitivity lacks the resonance that exists in the on-shell dark photon case; since at $m_{Z^{\prime}}\approx m_{V}$, the condition $m_{V}>2m_{\text{DM}}$ for the decay to be allowed, would not hold.

### v2

@@ 4 Direct Detection at DAMIC-M and PANDAX-4T
Here, we adopt the notation of Ref. Catena et al. (2020) where $m_{e}$ is the electron mass, $\boldsymbol{q}$ is the momentum transfer, $\boldsymbol{v}^{\perp}$ is the transverse DM-electron relative velocity while $\boldsymbol{S}_{e}^{s^{\prime}s}=\xi^{\dagger s^{\prime}}\boldsymbol{S}_{e}\xi^{s}$, $\boldsymbol{S}_{e}=\boldsymbol{\sigma}/2$ is the electron spin operator, $\boldsymbol{\sigma}=(\sigma_{1},\sigma_{2},\sigma_{3})$ is a three-dimensional vector whose components are the three Pauli matrices, $\xi^{s}$, $s=1,2$, is a two-component spinor, $\boldsymbol{S}_{DM}^{\lambda^{\prime}\lambda}=-i\boldsymbol{e}^{\prime}_{\lambda^{\prime}}\times\boldsymbol{e}_{\lambda}$, $\mathcal{S}^{\lambda^{\prime}\lambda}_{{\rm DM}\,ij}=\frac{1}{2}\left(e_{\lambda i}\,e^{\prime}_{\lambda^{\prime}j}+e_{\lambda j}\,e^{\prime}_{\lambda^{\prime}i}\right)$ and $e_{\lambda i}$ is the $i$-th component of the $\lambda$-th DM polarization vector. From Eq. LABEL:eq:MNR and the null results reported by DAMIC-M Aggarwal and others (2025) and PANDAX-4T Zhang and others (2025), we derive 90% C.L. exclusion limits in the $(m_{\rm DM},\sin\zeta)$ plane. For DAMIC-M, we perform this calculation by requiring that the probability of observing a total number of signal plus background events smaller than or equal to the actually observed one in the one, two and three electron channels is at least 0.1. In doing so, we use the expected background reported in Tab. 1 of Aggarwal and others (2025), and compute the rate of DM-induced electronic excitations in a silicon detector using the crystal response functions obtained in Catena et al. (2021). We assume the total detector efficiency given in Aggarwal and others (2025) and an exposure of 1.257 kg-year. For PANDAX-4T, we impose that, in each of the eight ionization-charge bins considered in Ref. Zhang and others (2025), the probability of observing a number of signal plus background events smaller than or equal to the actually observed one is at least 0.1. We extract the PANDAX-4T background model from Fig. 2 of Ref. Zhang and others (2025) and compute the differential electron recoil energy rate using the xenon response functions derived in Ref. Catena et al. (2020). We then convert this into an ionization rate by applying the detector response model introduced in Ref. Essig et al. (2012), the total detector efficiency reported in Fig. 1 of Ref. Zhang and others (2025) and employing an exposure of 1.04 ton-year. For both experiments, we assume Poisson statistics for the observed number of counts.
@@ 5 Sensitivity at Fixed-Target Experiments
In this section, we consider the most relevant experimental constraints and projections, including constraints from fixed-target experiments, colliders, direct detection, electroweak precision measurements, BBN, and from energy injection into the CMB.
@@
With the ongoing development of future fixed-target experiments, such as LDMX Appert and others (2025); Åkesson and others (2018), and the current competitive limits coming from NA64 Andreev and others (2025) and electron beam dump experiments Andreas et al. (2012), fixed-target experiments are capable of extensively probing sub-GeV DM for numerous theoretical scenarios Catena and Gray (2023); Catena et al. (2025); Berlin et al. (2020); Krnjaic and others (2022); Gori and others (2022); Cosme et al. (2021). Having a high energy beam of electrons (or protons in the case of SHiP Ahdida and others (2021) and other proton beam dumps) incident on a stationary target dense with nuclei, allows for a large number of potential DM events that can be detected through their missing energy or momentum (or through the direct scattering with a downstream detector in the case of proton beam dumps). However, for our scenario since $m_{Z^{\prime}}<m_{\text{DM}}$, production of DM proceeds through off-shell processes and is thus suppressed relative to its on-shell counterpart. The on-shell production of visible species, in particular $e^{+}e^{-}$ also occurs, acting as a complementary probe of this model. We consider the missing momentum search constraints from current NA64 null results and the sensitivity projections of the future experiment LDMX, in addition to the visible searches from past beam dump experiments E774 Bross et al. (1991), E141 Riordan and others (1987), Orsay Davier and Nguyen Ngoc (1989), E137 Bjorken et al. (1988), CHARM Gninenko (2012), KEK Konaka and others (1986), and NuCal Tsai et al. (2021).
@@
NA64’s 100 GeV electron beam incident on an active target composed of lead absorber layers, with $4.4\times 10^{11}$ electrons on target (EOT) so far, places competitive constraints on DM and dark sector particles that can be produced through these beam target interactions Andreev and others (2023, 2025). We consider both visible and invisible production, where for the latter the production of DM through dark bremsstrahlung and through the decay of vector mesons are each capable of delivering missing energy signatures.
@@
LDMX is a future missing-momentum experiment that will operate in two phases, where the second phase plans to have $10^{16}$ EOT with a beam energy of 8 GeV Åkesson and others (2023, 2018). With the ability to measure the transverse momentum, in addition to the energy, of the recoil electron after the interaction with the tungsten target, LDMX is subject to less background, projecting sensitivities reaching further than previous experiments in the on-shell case. In this study we explore the less studied and less sensitive off-shell / visible decay regime Berlin et al. (2019). Similarly to NA64, LDMX will search for signatures from visible decays, dark bremsstrahlung, and invisible vector meson decay Schuster et al. (2022).
@@ 5.1 Dark Bremsstrahlung
Beam electrons incident on a target could produce dark sector particles in a process similar to ordinary bremsstrahlung. In the usual dark photon kinetic mixing scenario where $m_{A^{\prime}}>2m_{\text{DM}}$ Berlin et al. (2020), the dark photon is produced on-shell and later decays to DM, therefore the event rate is independent of the DM properties and is only dependent on the dark photon. In our scenario however, there is the off-shell production of DM through a virtual $Z^{\prime}$ or ordinary photon, since the mass hierarchy is reversed such that $m_{\text{DM}}>m_{Z^{\prime}}$. Off-shell production of DM is suppressed compared to on-shell production Berlin et al. (2019), leading to less expected events. Dark bremsstrahlung processes that occur dominantly for our choice of model parameters, namely the emission of DM off the initial or final state electron, and dark Bethe-Heitler trident processes, are drawn in Fig. 4.
Figure 4: The dominant off-shell dark bremsstrahlung processes generating DM that occur at fixed-target experiments, where NN represents the target nucleus. Left: DM production diagrams off the initial or final state beam electron. Right: Bethe-Heitler trident diagrams for DM production.
Dark bremsstrahlung diagrams involving the new scalars, $\Phi_{D}^{0}$, $\Sigma_{D}^{0}$, and $\Sigma_{D}^{\pm}$, as intermediate states, become relevant as $\sin\beta$ approaches 0, however for the choice of parameters where $\cos\beta=0.7$, these diagrams contribute negligibly to the production of dark states. Simulations of dark bremsstrahlung events are performed in MadGraph Alwall et al. (2014), with UFO files Degrande et al. (2012) generated by our FeynRules model file Alloul et al. (2014). The DM and dark scalar vertices, in addition to the nucleus-photon vertex with form factor from Bjorken et al. (2009) [^3] The form factor from A18 and A19 of Bjorken et al. (2009), which is used as an approximate representation of the effects from the atomic/nuclear structure, has a typo in Eq. A19: the second term should not be squared., are included in our UFO files.
@@ 5.2 Dark Higgs-strahlung
Not only can the production of DM itself produce a missing momentum signature, but also the production of new scalars Acanfora and Kahlhoefer (2025); Duerr et al. (2017), since similarly to DM, dark scalars would not interact in the detector. Through the interaction vertices listed in Appendix C, dark scalar production occurs via dark Higgs-strahlung – analogously to normal dark bremsstrahlung. In our set up, the production of scalars $\Sigma_{D}^{\pm/0},\Phi_{D}^{0}$ through the process $e^{-}N\to e^{-}NS$, where $S$ is final states such as $\Sigma_{D}^{+(-)}X^{-(+)}$, $\Sigma_{D}^{+}\Sigma_{D}^{-}$, and $\Sigma_{D}^{0}/\Phi_{D}^{0}$, would look like a missing energy/momentum signature at an experiment like LDMX, similarly to a signature from DM production. The final states $\Sigma_{D}^{+(-)}X^{-(+)}$ make up the largest contribution of events. The processes with final states $\Phi_{D}^{0}Z^{\prime}$ and $\Sigma_{D}^{0}Z^{\prime}$ also occur, however there is only missing energy/momentum if the $Z^{\prime}$ decays visibly outside of the detector.
@@ 5.3 Invisible Vector Meson Decay
In addition to the dark bremsstrahlung process, DM can also be produced invisibly through vector meson decays generated in beam-target collisions – the so-called invisible vector meson decay Schuster et al. (2022). Bremsstrahlung photons are converted to vector mesons through exclusive photoproduction processes, which then decay to DM through a mediator that mixes with vector mesons. The three possible vector mediators in this model are the photon, $Z$ boson, and $Z^{\prime}$. The $Z$ boson is heavy compared to the vector mesons, resulting in a much smaller cross section than those of the other two mediators. There is a larger cross section through the ordinary photon, however not as high as through the $Z^{\prime}$ – which has a mass less than the DM but similar order of magnitude. Therefore, we consider the process $V\to Z^{\prime}\to X^{+}X^{-}$, where $V$ is a vector meson such as the $\rho$, $\omega$, $\phi$, or $J/\psi$, and the $Z^{\prime}$ is off-shell. Since the $Z^{\prime}$ couples to SM quarks through vector currents, the $Z^{\prime}$ mixes with the vector mesons with the form factors given in the fourth line of Table III of Schuster et al. (2022). Following the methodology of Schuster et al. (2022), and adapting to this model, we compute the predicted number of DM events from invisible vector meson decays for the experiments NA64 Andreev and others (2025) and LDMX Appert and others (2025); Åkesson and others (2018). The number of events from the decay of a vector meson $V$ is given by,
@@
where $N_{V}$ is the predicted number of vector mesons produced from beam target collisions and $Br(V\to X^{+}X^{-})\equiv\frac{\Gamma_{V\to X^{+}X^{-}}}{\Gamma_{V}}$ where the decay rate to DM is written in Appendix B and $\Gamma_{V}$ is measured experimentally and reported in PDG Tanabashi and others (2018). The sum runs over $V=\rho,\omega,\phi$ and in the case of NA64, additionally $J/\psi$ since the beam energy of 100 GeV is sufficiently high. We use the corresponding values found in Table II of Schuster et al. (2022) for $N_{V}$, calculated using a combination of experimental measurements and theory. We take the currently reported EOT at NA64 to be $4.4\times 10^{11}$, and thus scale-up the corresponding values in Table II of Schuster et al. (2022) to be $N_{\rho}=1.1\times 10^{7}$, $N_{\omega}=7.9\times 10^{5}$, $N_{\phi}=7.0\times 10^{5}$, and $N_{J/\psi}=9.7\times 10^{3}$. For LDMX, we take the corresponding values for Phase II reported in Table II of Schuster et al. (2022).
Figure 6: 90% C.L. constraints from current NA64 null results Andreev and others (2025) (in solid) and projections for LDMX Phase II Appert and others (2025) (in dashed), from invisible vector meson decay. The contribution from each vector meson is plotted in colour, and the total in black. $g_{D}$ is set to $10^{-2}$.
90% C.L. constraints and projections for invisible vector meson decays at NA64 and LDMX, respectively, for each specie of meson are plotted in Fig. 6 as a function of the model parameters. The reach of LDMX surpasses that of NA64 due to the increased statistics, providing a significant improvement to the sensitivity landscape from the data that will come from LDMX. Interestingly, since the $V\to X^{+}X^{-}$ decay rate is proportional to $1/m_{\text{DM}}^{4}$, the sensitivity increases as $m_{\text{DM}}$ decreases – resulting in much more sensitivity for lower masses compared to the on-shell dark photon studied in Schuster et al. (2022), which features a sensitivity that is independent of $m_{\text{DM}}$. Due to the inverse mass hierarchy, where $m_{\text{DM}}>m_{Z^{\prime}}$, the sensitivity lacks the resonance that exists in the on-shell dark photon case; since at $m_{Z^{\prime}}\approx m_{V}$, the condition $m_{V}>2m_{\text{DM}}$ for the decay to be allowed, would not hold.